# Ground What You See: Hallucination-Resistant MLLMs via Caption Feedback, Diversity-Aware Sampling, and Conflict Regularization

### v1

@@ Abstract
Code — https://github.com/omniAI-Lab/Ground-What-You-See.git
@@ Introduction
Baseline. To evaluate our method, we compare it against a range of representative baselines. GPT-4o (Hurst et al. 2024) and Gemini-1.5-Pro (Team et al. 2024) serve as strong closed-source performance references. Open-source video models, including LLaVA-OneVision-7B (Li et al. 2024) and VILA-1.5-8B, are relevant for temporal visual understanding. General-purpose vision-language foundations such as MiniCPM-V2.6-8B (Yao et al. 2024), InternVL2.5-8B (Chen et al. 2024), and LLaMA3.2-11B (Grattafiori et al. 2024) are used as base models. Reasoning-focused baselines like R1-OneVision-7B (Yang et al. 2025), R1-VL (Zhang et al. 2025), Vision-R1-7B (Huang et al. 2025), and Video-R1-7B (Feng et al. 2025) incorporate structured prompting or alignment for improved multi-step reasoning. Finally, we include our models: the base Qwen-VL-2.5-7B (Bai et al. 2025), its SFT and GRPO variants, and our enhanced version (Ours), to analyze the impact of visual grounding and contrastive regularization. All baselines are compared under consistent settings to assess performance in mitigating hallucinations and enhancing reasoning robustness.
@@ Conclusion
Limitations and Future Work. While contrastive regularization enhances robustness, it introduces mild sensitivity to hyperparameters such as the carefully tuned optimal NTK similarity threshold. Future work will extend our analysis to track reasoning drift over longer token spans, integrate lightweight captioning modules into frozen visual backbones for broader applicability, and adapt our methods to instruction-following tasks in real-world video applications where grounding and reasoning reliability are paramount.
@@ Acknowledgments
Appendix A Proof
Proof of Eq. (2)
Although GRPO includes clipped advantages and importance sampling ratio corrections, it is fundamentally a gradient-based policy optimization method. In particular, the ratio correction term—defined as the likelihood ratio between the target policy and a fixed reference policy—acts as a constant denominator during optimization, while only the numerator depends on the trainable model. Therefore, for gradient analysis, we may safely approximate the objective as a reward-weighted softmax and omit both clipping and reweighting terms. Specifically, at each decoding step , the GRPO loss for the sampled token can be written as:
where denotes the logit of token , and is the scalar advantage signal associated with the sampled token .
We now compute the gradient of this loss with respect to the logits , treating as a constant scalar.
Case 1: (sampled token)
Case 2: (unsampled token)
In summary, the gradient of the GRPO loss with respect to the logits can be written as:
This result highlights how GRPO updates both sampled and unsampled tokens in a reward-aware manner: increasing the logit of when , and softly decaying competing tokens based on their prior probabilities.
Proof of Eq. (3)
Under GRPO, the loss for training example is defined as:
Ignore the clipped operator, the gradient of the loss with respect to logits is:
Suppose we want to observe the model’s prediction on an “observing example” . Starting from the first-order Taylor expansion, we write:
Assuming the model updates its parameters using SGD calculated by an “updating example” , we rearrange the above to get:
To evaluate the leading term, we expand using the chain rule:
Recalling the definitions:
For the higher-order term, using the same update:
and assuming is bounded or clipped in practice, we have:
Substituting into the above, we obtain:
Appendix B Training Details
We fine-tune a large-scale video-language model using the GRPO framework. Video inputs are preprocessed into 16-frame clips, uniformly sampled from the raw videos. A unified tokenizer is employed to process both visual tags and natural language text. The fine-tuning is conducted on 8 NVIDIA A800 GPUs with 80GB memory each. For each GRPO iteration, we sample candidate responses per input prompt (video-question pair) using a generation temperature of 1.2. The rollout batch size is set to 1, We use the Adam optimizer with separate learning rates of for the actor and for the critic. To prevent overfitting to sampled rewards, a KL-penalty coefficient of 0.04 is applied to constrain the updated policy within a trust region defined by the reference model. The entire fine-tuning process runs for 3,000 steps, with evaluation conducted every 100 steps using a fixed set of held-out prompts.
Our model underwent two stages of training: the cold-start stage and the reinforcement learning stage. For training with video data, we employed uniform sampling, selecting 16 frames as input for video training.
Cold-start Stage: In the Cold-start stage, we curated and modified the LLaVA-CoT-100k dataset to align with our proposed output schema. Specifically, we leveraged DeepSeek to adjust the SUMMARY fields within the dataset to match the structure of our defined planning output stage. Based on the resulting dataset, we conducted supervised fine-tuning on the Qwen2.5-VL-7B-Instruct model for 2 epochs. This process yielded a cold-start version of the model, referred to as Qwen2.5-VL-7B-SFT.
Reinforcement Learning Stage: In the reinforcement learning stage, we selected a subset of 40k samples from the Video-R1 dataset to form our initial training set. The selection procedure is described in detail in Appendix D. To refine the dataset further, we applied our proposed sample selection method to identify 20k high-quality samples, which were then used as the final training set for reinforcement learning. We conducted the reinforcement learning process on the Qwen2.5-VL-7B-SFT model obtained from the cold-start phase. Training was performed on 8 A800 (80GB) GPUs over the span of approximately three days, completing a total of 3000 steps, which resulted in the Qwen2.5-VL-7B-Ours model.
Appendix C More Examples
Appendix D Complete list of training data
We selected 38k data from the Video-R1 open source dataset as our training set. The specific composition of the selected data is as follows:
| Dataset | Name | Size | |
| Image | Chart | TabMWP | 2k |
| MapQA | 1k | ||
| FigureQA | 1k | ||
| DVQA | 1.2k | ||
| PlotQA | 0.8k | ||
| A-OKVQA | 2k | ||
| General | Visual7W | 1k | |
| IconQA | 2k | ||
| ArxivQA | 2k | ||
| Knowledge | ScienceQA | 1k | |
| AI2D | 2k | ||
| EXAMS-V | 1k | ||
| Math | Multimath-300k | 3k | |
| Spatial | OpenSpaces | 1k | |
| Spacellava | 1k | ||
| Video | LLaVA-Video-178K | 15k | |
| STAR | 3k | ||

### v2

@@ Abstract
Code — https://github.com/ZJU-OmniAI/OMNEX-VL
@@ Introduction
Baseline. To evaluate our method, we compare it against a range of representative baselines. GPT-4o (Hurst et al. 2024) and Gemini-1.5-Pro (Team et al. 2024) serve as strong closed-source performance references. Open-source video models, including LLaVA-OneVision-7B (Li et al. 2024) and VILA-1.5-8B, are relevant for temporal visual understanding. General-purpose vision-language foundations such as MiniCPM-V2.6-8B (Yao et al. 2024), InternVL2.5-8B (Chen et al. 2024), and LLaMA3.2-11B (Grattafiori et al. 2024) are used as base models. Reasoning-focused baselines like R1-OneVision-7B (Yang et al. 2025), R1-VL (Zhang et al. 2025), Vision-R1-7B (Huang et al. 2025), and Video-R1-7B (Feng et al. 2025) incorporate structured prompting or alignment for improved multi-step reasoning. Finally, we include our models: the base Qwen-VL-2.5-7B (Bai et al. 2025), its SFT and GRPO variants, and our enhanced version (Ours), to analyze the impact of visual grounding and contrastive regularization.